OpenStudy (anonymous):

Progession The sixth term of an arithmetic progession is 265 and the sum of the first 5 terms is 1445. Find the minimun value of n so that the sum of the first n terms is negative.

OpenStudy (anonymous):

@ParthKohli

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

Do you know the formulae? you can form 2 equations with a1 and d using the given info....

OpenStudy (anonymous):

yes \[265=a+5d-First~eqn\]\[289=a+2d-Second~eqn\]

OpenStudy (anonymous):

\[First~term,a=305,common~difference,d=-8\] is it correct?

hartnn (hartnn):

yes! good work :)

OpenStudy (anonymous):

but i'm not sure what do next...i'm stuck at here :(

OpenStudy (anonymous):

okay,what should we do next?

hartnn (hartnn):

lets setup the sum inequality we want the sum of 'n' terms to be negative, \((n/2) (2a + (n-1)d) < 0 \) plug in a and d values in this

OpenStudy (anonymous):

\[\frac{ n }{ 2 }(2(305)-(n-1)(-8))\]

hartnn (hartnn):

yup, keep on simplifying it

OpenStudy (anonymous):

\[-4n^2+309n<0\]

hartnn (hartnn):

Remember : n cannot be negative

OpenStudy (anonymous):

4n^2-309n>0

hartnn (hartnn):

how did u get 309 ?

hartnn (hartnn):

n (305 - (-4)*(n-1)) < 0 n < 0 or (305 - (-4)*(n-1) < 0 discarding n<0 305 +4(n-1) < 0 301 + 4n < 0 ....

hartnn (hartnn):

got that?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

n>-305/4

hartnn (hartnn):

the formula has a + in between

hartnn (hartnn):

(305 + (-4)*(n-1) < 0

hartnn (hartnn):

(305 + 4 - 4n) < 0 309 -4n < 0 4n > 309/4

OpenStudy (anonymous):

is n>77.25

hartnn (hartnn):

** n > 309/4 n > 77.25

hartnn (hartnn):

yes

OpenStudy (anonymous):

so answer is n=78

hartnn (hartnn):

\(\huge \checkmark \)

OpenStudy (anonymous):

Thnx @hartnn :)

hartnn (hartnn):

welcome ^_^